# Value Flows

### v1

@@ F.2 Methods for Comparison
Implicit Q-Learning is an offline RL method that uses expectile regression to learn the best action within the support of the behavioral dataset. Following prior work (Park et al., 2025c), we perform hyperparameter search over the AWR inverse temperatures $\alpha$ in $\{0.3,1,3,10\}$ for each environment (Table 4).
@@
IFQL is a variant of Implicit Diffusion Q-Learning (IDQL; (Hansen-Estruch et al., 2023a)) with flow policy. IFQL uses the flow policy to propose candidates and uses rejection sampling to select actions as in Value Flows. We select the number of candidates in rejection sampling NN from $\{32,64,128\}$ (Table 4).
@@
C51 is a distributional RL method that discretizes the return distribution into a fixed number of bins and uses cross-entropy loss to update these categorical distributions. We tune the number of atoms $N$ in $\{51,101\}$ (Table 4).
@@
IQN is a distributional RL method that approximates the return distribution by predicting quantile values at randomly sampled quantile fractions. We select the temperature $\kappa$ in the quantile regression loss from $\{0.7,0.8,0.9,0.95\}$ (Table 4).
@@
CODAC combines the IQN critic loss with conservative constraints. Similar to Park et al. (2025c), we use a one-step flow policy to perform DDPG-style policy extraction. We fixed the conservative penalty coefficient to $0.1$. We sweep the temperature $\kappa$ in the quantile regression loss over $\{0.7,0.8,0.9,0.95\}$ and sweep the BC coefficient $\alpha_{1}$ over $\{100,300,1000,3000,10000,30000\}$ (Table 4).
F.3 Hyperparameters
Table 4: Domain specific hyperparameters for Value Flows and baselines. The complete descriptions of each hyperparameter can be found in Appendix F.2. Following prior work Park et al. (2025c), we tune these hyperparameters for each domain from OGBench benchmarks on the “*” task. “-” indicates the results were taken from prior work (Park et al., 2025c) (or do not exist). The discount factor $\gamma$ is shared for all methods in each domain.
Discount IQL ReBRAC FBRAC FQL IFQL C51 IQN CODAC Value Flows
Domain or task $\gamma$ $\alpha$ $\alpha_{1}$ $\alpha_{2}$ $\alpha$ $\alpha$ NN NN $\kappa$ $\kappa$ $\alpha$ $\lambda$ $\tau$
cube-double-play $0.995$ $0.3$ $0.1$ 0 $100$ $100$ $32$ $101$ $0.9$ $0.95$ $300$ 11 33
cube-triple-play $0.995$ $10$ $0.03$ 0 $100$ $300$ $32$ $51$ $0.8$ $0.95$ $1000$ 33 $0.03$
puzzle-3x3-play $0.99$ $10$ $0.3$ $0.001$ $100$ $1000$ $32$ $101$ $0.8$ $0.95$ $100$ $0.5$ $0.3$
puzzle-4x4-play $0.99$ 33 $0.3$ $0.01$ $300$ $1000$ $32$ $101$ $0.95$ $0.95$ $1000$ 33 $100$
scene-play $0.99$ $10$ $0.1$ $0.001$ $100$ $300$ $32$ $51$ $0.95$ $0.95$ $100$ 11 $0.3$
visual-antmaze-medium-navigate $0.99$ 11 $0.01$ $0.003$ $100$ $100$ $32$ - $0.9$ - - $0.3$ $0.03$
visual-antmaze-teleport-navigate $0.99$ 11 $0.01$ $0.003$ $100$ $100$ $32$ - 0.8 - - $0.3$ $0.03$
visual-cube-double-play $0.995$ $0.3$ $0.1$ 0 $100$ $100$ $32$ - $0.9$ - - 11 $0.3$
visual-puzzle-3x3-play $0.99$ $10$ $0.3$ $0.01$ $100$ $300$ $32$ - $0.8$ - - $0.3$ $0.3$
visual-scene-play $0.99$ $10$ $0.1$ $0.003$ $100$ $100$ $32$ - $0.95$ - - 11 $0.3$
pen-human $0.99$ - - - $30000$ $10000$ $32$ $51$ $0.8$ $0.8$ $10000$ 33 11
pen-cloned $0.99$ - - - $10000$ $10000$ $32$ $51$ $0.8$ $0.8$ $10000$ 33 11
pen-expert $0.99$ - - - $30000$ $3000$ $32$ $51$ $0.8$ $0.8$ $10000$ 33 $0.01$
door-human $0.99$ - - - $30000$ $30000$ $32$ $101$ $0.9$ $0.9$ $10000$ 33 $0.01$
door-cloned $0.99$ - - - $10000$ $30000$ $32$ $51$ $0.9$ $0.9$ $30000$ $10$ $0.3$
door-expert $0.99$ - - - $30000$ $30000$ $32$ $51$ $0.9$ $0.9$ $10000$ $10$ $0.3$
hammer-human $0.99$ - - - $30000$ $30000$ $128$ $51$ $0.7$ $0.8$ $30000$ 33 $0.3$
hammer-cloned $0.99$ - - - $10000$ $10000$ $32$ $51$ $0.7$ $0.8$ $10000$ 33 $0.3$
hammer-expert $0.99$ - - - $30000$ $30000$ $32$ $51$ $0.9$ $0.8$ $10000$ $10$ 11
relocate-human $0.99$ - - - $30000$ $10000$ $32$ $101$ $0.9$ $0.9$ $30000$ $10$ $0.01$
relocate-cloned $0.99$ - - - $3000$ $30000$ $32$ $51$ $0.9$ $0.9$ $30000$ 33 $0.01$
relocate-expert $0.99$ - - - $30000$ $30000$ $32$ $101$ $0.9$ $0.9$ $10000$ 33 $0.1$

### v2

@@ E.2 Methods for Comparison
Implicit Q-Learning is an offline RL method that uses expectile regression to learn the best action within the support of the behavioral dataset. Following prior work (Park et al., 2025c), we learn a Gaussian policy via AWR and perform hyperparameter search over the AWR inverse temperatures $\alpha$ in $\{0.3,1,3,10\}$ for each environment (Table 6).
@@
IFQL is a variant of Implicit Diffusion Q-Learning (IDQL; (Hansen-Estruch et al., 2023a)) with flow policy. IFQL uses the flow policy to propose candidates and uses rejection sampling to select actions as in Value Flows. We select the number of candidates in rejection sampling $N$ from $\{32,64,128\}$ (Table 6).
@@
C51 is a distributional RL method that discretizes the return distribution into a fixed number of bins and uses cross-entropy loss to update these categorical distributions. We tune the number of atoms $N$ in $\{51,101\}$ (Table 6). Because the original C51 algorithm aims to solve the Atari game with a discrete action space, it greedily selects actions that maximize the Q-value estimated by the distributional returns (no explicit policy). However, since our evaluation benchmarks (Appendix E.1) have continuous action spaces, we choose to learn an explicit flow BC policy and use rejection sampling to select actions (Appendix C.2), resembling the greedy action selection in a discrete action space. Using flow-based rejection sampling for C51 not only matches the policy extraction method of Value Flows, but also prevents sampling out-of-distribution actions (Chen et al., 2023).
@@
IQN is a distributional RL method that approximates the return distribution by predicting quantile values at randomly sampled quantile fractions. We select the temperature $\kappa$ in the quantile regression loss from $\{0.7,0.8,0.9,0.95\}$ (Table 6). Similar to the adaptation in C51, we also learn an explicit flow BC policy and use rejection sampling (Appendix C.2) to select actions that maximize the Q estimated by IQN.
@@
CODAC combines the IQN critic loss with conservative constraints. Similar to Park et al. (2025c), we use a one-step flow policy to perform DDPG-style policy extraction (Appendix C.2). We fixed the conservative penalty coefficient to $0.1$. We sweep the temperature $\kappa$ in the quantile regression loss over $\{0.7,0.8,0.9,0.95\}$ and sweep the BC coefficient $\alpha_{1}$ over $\{100,300,1000,3000,10000,30000\}$ (Table 6).
Value Flows.
Our method, Value Flows, uses flexible flow-based models to estimate the full distribution of future returns and to identify epistemic uncertainty across different states using the learned generative flow models. We describe the components of the practical Value Flows algorithm in Appendix C.1. In our experiments, we sweep the BCFM regularization coefficient $\lambda\in\{0.3,0.5,1,3,10\}$ and the confidence weight temperature $\tau\in\{0.01,0.03,0.1,0.3,1,3\}$. In addition, the flow-based distillation strength $\alpha$ is the key hyperparameter for our offline-to-online experiments. We select the distillation strength $\alpha$ from $\{10,30,100,300\}$. See Table 6 for the specific values of the hyperparameter for each domain or task.
E.3 Hyperparameters
@@
Table 6: Domain specific hyperparameters for Value Flows and baselines. The complete descriptions of each hyperparameter can be found in Appendix E.2. Following prior work Park et al. (2025c), we tune these hyperparameters for each domain from OGBench benchmarks on the default (“*”) task in Table 3. “-” indicates the values do not exist. [^6] The hyperparameters for antmaze-large-navigate and humanoidmaze-medium-navigate are used for offline-to-online experiments.
Discount IQL ReBRAC FBRAC IFQL FQL C51 IQN CODAC Value Flows
Domain or task $\gamma$ $\alpha$ $\alpha_{1}$ $\alpha_{2}$ $\alpha$ NN $\alpha$ NN $\kappa$ $\kappa$ $\alpha$ $\lambda$ $\tau$ $\alpha$
antmaze-large-navigate $0.99$ $10$ - - - $32$ $10$ - $0.7$ - - $10$ $0.01$ $30$
humanoidmaze-medium-navigate $0.995$ $10$ - - - $32$ $30$ - $0.7$ - - 33 $0.3$ $100$
cube-double-play $0.995$ $0.3$ $0.1$ 0 $100$ $32$ $100$ $101$ $0.9$ $0.95$ $300$ 11 33 $300$
cube-triple-play $0.995$ $10$ $0.03$ 0 $100$ $32$ $300$ $51$ $0.8$ $0.95$ $1000$ 33 $0.03$ $300$
puzzle-3x3-play $0.99$ $10$ $0.3$ $0.001$ $100$ $32$ $1000$ $101$ $0.8$ $0.95$ $100$ $0.5$ $0.3$ -
puzzle-4x4-play $0.99$ 33 $0.3$ $0.01$ $300$ $32$ $1000$ $101$ $0.95$ $0.95$ $1000$ 33 $100$ $300$
scene-play $0.99$ $10$ $0.1$ $0.001$ $100$ $32$ $300$ $51$ $0.95$ $0.95$ $100$ 11 $0.3$ $300$
visual-antmaze-medium-navigate $0.99$ 11 $0.01$ $0.003$ $100$ $32$ $100$ - $0.9$ - - $0.3$ $0.03$ -
visual-antmaze-teleport-navigate $0.99$ 11 $0.01$ $0.003$ $100$ $32$ $100$ - 0.8 - - $0.3$ $0.03$ -
visual-cube-double-play $0.995$ $0.3$ $0.1$ 0 $100$ $32$ $100$ - $0.9$ - - 11 $0.3$ -
visual-puzzle-3x3-play $0.99$ $10$ $0.3$ $0.01$ $100$ $32$ $300$ - $0.8$ - - $0.3$ $0.3$ -
visual-scene-play $0.99$ $10$ $0.1$ $0.003$ $100$ $32$ $100$ - $0.95$ - - 11 $0.3$ -
pen-human $0.99$ - - - $30000$ $32$ $10000$ $51$ $0.8$ $0.8$ $10000$ 33 11 -
pen-cloned $0.99$ - - - $10000$ $32$ $10000$ $51$ $0.8$ $0.8$ $10000$ 33 11 -
pen-expert $0.99$ - - - $30000$ $32$ $3000$ $51$ $0.8$ $0.8$ $10000$ 33 $0.01$ -
door-human $0.99$ - - - $30000$ $32$ $30000$ $101$ $0.9$ $0.9$ $10000$ 33 $0.01$ -
door-cloned $0.99$ - - - $10000$ $128$ $30000$ $51$ $0.9$ $0.9$ $30000$ $10$ $0.3$ -
door-expert $0.99$ - - - $30000$ $32$ $30000$ $51$ $0.9$ $0.9$ $10000$ $10$ $0.3$ -
hammer-human $0.99$ - - - $30000$ $32$ $30000$ $51$ $0.7$ $0.8$ $30000$ 33 $0.3$ -
hammer-cloned $0.99$ - - - $10000$ $32$ $10000$ $51$ $0.7$ $0.8$ $10000$ 33 $0.3$ -
hammer-expert $0.99$ - - - $30000$ $32$ $30000$ $51$ $0.9$ $0.8$ $10000$ $10$ 11 -
relocate-human $0.99$ - - - $30000$ $128$ $10000$ $101$ $0.9$ $0.9$ $30000$ $10$ $0.01$ -
relocate-cloned $0.99$ - - - $3000$ $32$ $30000$ $51$ $0.9$ $0.9$ $30000$ 33 $0.01$ -
relocate-expert $0.99$ - - - $30000$ $32$ $30000$ $101$ $0.9$ $0.9$ $10000$ 33 $0.1$ -
We include common hyperparameters for Value Flows and baselines in Table 5 and domain-specific hyperparameters for different methods in Table 6.